OpenStudy (anonymous):

Solve for g 5g-2h=3g+1

OpenStudy (anonymous):

bring all like terms on i side: 5g-3g=2h+1 simplify: 2g=2h+1 divide 2 to get g by itself: g=h+1/2

OpenStudy (hackergod):

first, you have to put all the alike terms on one side. 5g-3g=2h+1, then you simplify. 5-3= 2g .... 2g=2h+1.. divide by 2 to get g by itself... you get the answe g=h+1/2

OpenStudy (anonymous):

its says its wrong